# On the sample complexity of estimation in logistic regression

### v1

@@ 3 Upper bounds on the sample complexity
There is an absolute constant $C>0$ and an estimator ${\hat{\theta}}$ such that, for any $\theta^{\star}\in S^{d-1}$ and any $\epsilon,\delta\in\lparen 0,1\rparen$ ), we have
$\Pr\lparen\lVert{\hat{\theta}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert\leq\epsilon\rparen\geq 1-\delta,$
where the probability is with respect to the data with distribution determined by $\theta^{\star}$, provided that
@@
$n\geq C\min\left\{\left\lparen\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\beta\varepsilon^{2}}+\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right\rparen,\frac{d+\log(1/\delta)}{\beta^{2}\varepsilon^{2}}\right\}.$
The estimator referenced in Theorem 3 is based on empirical risk minimization. We describe this estimator and its properties below. Fix a “ground truth” parameter $\theta^{\star}\in S^{d-1}$ determining the distribution of $(\mathbf{x},\mathbf{y})$, and note that
The empirical risk minimizer ${\hat{\theta}_{\operatorname{ERM}}}={\hat{\theta}_{\operatorname{ERM}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})$ is the $\theta\in S^{d-1}$ that minimizes
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta\leq 0\}.$
@@
(If the minimizer not unique, we choose one arbitrarily.) The following is a standard performance guarantee for empirical risk minimization, combined with the fact that the VC dimension of the class of homogeneous linear classifiers is $d$ [^3] Note that Lemma 5 is true under every distribution for $(\mathbf{x},\mathbf{y})$; it does not rely on specific properties of our model. In the special case of where $\mathbf{x}$ is the $d$-dimensional standard normal (or any other spherically symmetric distribution) and $\beta=\infty$ (which implies $\operatorname{opt}=0$), the $\log(1/\varepsilon)$ can be removed (Long, 1995). So, in this case, the sample size requirement is $C(d+\log(1/\delta))/\varepsilon$.
Lemma 5 (Vapnik and Chervonenkis, 1971; Talagrand, 1994).
@@
$n\geq C\min\left\{\left\lparen\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}\operatorname{opt}+\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right\rparen,\frac{d+\log(1/\delta)}{\varepsilon^{2}}\right\},$
@@
The following lemma is used to relate $\lVert{\hat{\theta}_{\operatorname{ERM}}}-\theta^{\star}\rVert$ to $\operatorname{err}_{\theta^{\star}}({\hat{\theta}_{\operatorname{ERM}}})-\operatorname{err}_{\theta^{\star}}(\theta^{\star})$.
@@
If $\theta\in S^{d-1}$ satisfies
$\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\varepsilon,$
then
$\lVert\theta-\theta^{\star}\rVert\leq\frac{\pi\varepsilon}{1-2\operatorname{err}_{\theta^{\star}}(\theta^{\star})}.$
Proof.
The distribution of $\mathbf{y}$ depends on $\mathbf{x}$ only through $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$. Therefore, since $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta$ and $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$ are each standard normal random variables with correlation $\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$, Lemma 11 implies
$\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\mid\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}>0)=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}=1-\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}>0\mid\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0).$
Therefore, the error rate of $\theta$ is
$\operatorname{err}_{\theta^{\star}}(\theta)=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\left\lparen 1-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen+\left\lparen 1-\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\right\rparen\operatorname{err}_{\theta^{\star}}(\theta^{\star}),$
and hence
$\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\left\lparen 1-2\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen\geq\frac{\lVert\theta-\theta^{\star}\rVert}{\pi}\left\lparen 1-2\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen$ ) )
where the inequality uses $\arccos(\rho)\geq\sqrt{2(1-\rho)}$. Combining this inequality with the hypothesis that $\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\varepsilon$ completes the proof. ∎
Finally, to deduce upper bounds on the sample complexity for achieving $\lVert{\hat{\theta}_{\operatorname{ERM}}}-\theta^{\star}\rVert\leq\epsilon$, we combine Lemma 5 and Lemma 6 with the bound on $\operatorname{opt}=\operatorname{err}_{\theta^{\star}}(\theta^{\star})$ from Lemma 4.

### v3

@@ 3.2 Moderate and low temperatures
Assume $\beta\geq 1+c$ for some positive constant $c>0$. Fix any $\epsilon,\delta\in\lparen 0,1\rparen$ ), and suppose
@@
$n\geq C\left\lparen\frac{d\log(d/\epsilon)+\log(1/\delta)}{\beta\epsilon^{2}}+\frac{d\log(d/\epsilon)+\log(1/\delta)}{\epsilon}\right\rparen$ )
where $C>0$ is an absolute constant. Then with probability at least $1-\delta$, we have
$\lVert{\hat{\theta}_{\operatorname{relu}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert\leq\epsilon.$
The main technical work in the proof of Theorem 3 is understanding the concentration properties of random variables of the form
@@
$\boldsymbol{\delta}_{\theta}:=[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+}-[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}]_{+},\quad\theta\in S^{d-1}.$
This was studied by Kuchelmeister and van de Geer (2023) in the case of the probit regression model. We obtain the necessary moment bounds for the logistic regression model. The following lemma (proved in Appendix B) gives the required bounds.
Lemma 5.
For any $\theta\in S^{d-1}$, and any integer $q\geq 2$,
$\operatorname{\mathbb{E}}\left[\lvert\boldsymbol{\delta}_{\theta}\rvert^{q}\right]\leq\frac{q!v}{2}b^{q-2}$
where
$b=C\lVert\theta-\theta^{\star}\rVert\quad\text{and}\quad v\leq C\lVert\theta-\theta^{\star}\rVert^{2}\left\lparen\frac{1}{\beta}+\lVert\theta-\theta^{\star}\rVert\right\rparen$ )
and $C>0$ is an absolute constant.
Given Lemma 5, the rest of the analysis is mostly standard. Lemma 5 and Bernstein’s inequality implies that each random variable in the empirical process above is concentrated around its expectation.
Lemma 6.
For any $\theta\in S^{d-1}$, define $\boldsymbol{\delta}_{\theta}^{i}:=[-\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+}-[-\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}]_{+}$ for all $i=1,\dotsc,n$. For any $\delta\in\lparen 0,1\rparen$ ), with probability at least $1-\delta$, we have
$\operatorname{\mathbb{E}}\left[\boldsymbol{\delta}_{\theta}\right]-\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{\delta}_{\theta}^{i}\leq C\lVert\theta-\theta^{\star}\rVert\sqrt{\left\lparen\frac{1}{\beta}+\lVert\theta-\theta^{\star}\rVert\right\rparen\frac{\log(1/\delta)}{n}}+C\lVert\theta-\theta^{\star}\rVert\frac{\log(1/\delta)}{n},$
where $C>0$ is an absolute constant.
To bound all random variables in the stochastic process simultaneously, we use a covering argument, again following Kuchelmeister and van de Geer (2023).
Lemma 7 (Kuchelmeister and van de Geer, 2023, Lemma 5.1.2).
For any $\varepsilon_{0}\in\lparen 0,1\rparen$ ) and any $A\subseteq S^{d-1}$, let $T(A,\varepsilon_{0})$ be an $\varepsilon_{0}$-cover of $A$ with respect to $\ell_{2}$ distance. Define $\boldsymbol{\delta}_{\theta}^{i}:=[-\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+}-[-\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}]_{+}$ for all $i=1,\dotsc,n$ and all $\theta\in A$. For any $\delta\in\lparen 0,1\rparen$ ), with probability at least $1-\delta$, we have
$\sup_{\theta\in A}\left\{\operatorname{\mathbb{E}}\left[\boldsymbol{\delta}_{\theta}\right]-\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{\delta}_{\theta}^{i}\right\}\leq\max_{\theta\in T(A,\varepsilon_{0})}\left\{\operatorname{\mathbb{E}}\left[\boldsymbol{\delta}_{\theta}\right]-\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{\delta}_{\theta}^{i}\right\}+\varepsilon_{0}\left\lparen\sqrt{\frac{2\ln(1/\delta)}{n}}+2\sqrt{d}\right\rparen.$ ) .
To relate $\operatorname{\mathbb{E}}[\boldsymbol{\delta}_{\theta}]$ to $\lVert\theta-\theta^{\star}\rVert$, we use the following lemma.
Lemma 8.
For any $\theta\in S^{d-1}$, we have
$\operatorname{\mathbb{E}}\left[\boldsymbol{\delta}_{\theta}\right]\geq\frac{1}{8}\sqrt{\frac{2}{\pi}}\left\lparen 1-\frac{1}{\beta^{2}}\right\rparen\lVert\theta-\theta^{\star}\rVert^{2}.$
Proof.
The expected KL divergence of $\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta))$ from $\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}))$ is the expected excess logistic loss of $w:=\beta\theta$ compared to $w^{\star}:=\beta\theta^{\star}$:
$\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})))\right]$ $=\operatorname{\mathbb{E}}\left[\ln(1+e^{-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w})\right]-\operatorname{\mathbb{E}}\left[\ln(1+e^{-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w^{\star}})\right]$
(recalling that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}=x$ is $\operatorname{Bern}(g^{\prime}(x^{\scriptscriptstyle{\mathsf{T}}}w^{\star}))$). Furthermore, the logistic loss $w\mapsto\ln(1+e^{-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w})$ decomposes into the sum of a label-dependent part and a label-independent part:
$\ln(1+e^{-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w})=\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w\right]_{+}+\ln(1+e^{-\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w\rvert}).$
(This decomposition is the same as that from Kuchelmeister and van de Geer (2023), and a similar decomposition was used by Bach (2010).) We can thus write the excess expected logistic loss as
$\operatorname{\mathbb{E}}\left[\ln(1+e^{-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w})\right]-\operatorname{\mathbb{E}}\left[\ln(1+e^{-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w^{\star}})\right]$ $=\operatorname{\mathbb{E}}\left[\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w\right]_{+}\right]+\operatorname{\mathbb{E}}\left[\ln(1+e^{-\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w\rvert})\right]$
$\qquad-\operatorname{\mathbb{E}}\left[\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w^{\star}\right]_{+}\right]-\operatorname{\mathbb{E}}\left[\ln(1+e^{-\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w^{\star}\rvert})\right]$
$=\operatorname{\mathbb{E}}\left[\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w\right]_{+}-\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w^{\star}\right]_{+}\right]$
since $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w$ and $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}w^{\star}$ have the same distribution. Therefore, we have
$\operatorname{\mathbb{E}}\left[\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta\right]_{+}-\left[-\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\right]_{+}\right]=\frac{1}{\beta}\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})))\right].$
The claim now follows by applying Lemma 15 (see Appendix A). ∎
Proof of Theorem 3.
Fix some $\varepsilon_{0},r_{0}\in\lparen 0,1\rparen$ ) and let $r_{j}=2^{j}r_{0}$ for $1\leq j\leq J:=\lceil\log_{2}(2/r_{0})\rceil$. Define $A_{0}=\{\theta\in S^{d-1}:\lVert\theta-\theta^{\star}\rVert\leq r_{0}\}$ and $A_{j}=\{\theta\in S^{d-1}:r_{j-1}<\lVert\theta-\theta^{\star}\rVert\leq r_{j}\}$ for all $j\geq 1$. Let $T_{j}=T(A_{j},\varepsilon_{0}(r_{j}/2)^{3/2})$ be an $\varepsilon_{0}(r_{j}/2)^{3/2}$-cover of $A_{j}$ with respect to $\ell_{2}$ distance of cardinality at most $(2(2/r_{j})^{3/2}\varepsilon_{0}^{-1})^{d-1}$ (cf. Lemma 3). Fix $\delta_{0}\in\lparen 0,1/(J+\sum_{j=0}^{J}\lvert T_{j}\rvert)\rparen$ ), and apply Lemma 7 to each $A_{j}$ for $1\leq j\leq J$, and Lemma 6 to each $\theta\in T$. By a union bound, with probability at least $1-(J+\sum_{j=0}^{J}\lvert T_{j}\rvert)\delta_{0}$, we have for all $\theta\in S^{d-1}$,
$\operatorname{\mathbb{E}}\left[\boldsymbol{\delta}_{\theta}\right]$ $\leq\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{\delta}_{\theta}^{i}+2C^{\prime}\Delta_{r_{0}}(\theta)\sqrt{\left\lparen\frac{1}{\beta}+\Delta_{r_{0}}(\theta)\right\rparen\frac{\log(1/\delta_{0})}{n}}+2C^{\prime}\Delta_{r_{0}}(\theta)\frac{\log(1/\delta_{0})}{n}$
$\qquad+\varepsilon_{0}\Delta_{r_{0}}(\theta)^{3/2}\left\lparen\sqrt{\frac{2\ln(1/\delta_{0})}{n}}+2\sqrt{d}\right\rparen$ )
where $\Delta_{r_{0}}(\theta):=\max\{r_{0},\lVert\theta-\theta^{\star}\rVert\}$, $C^{\prime}>0$ is some absolute constant, and $\boldsymbol{\delta}_{\theta}^{i}$ are as in Lemma 6. We condition on the event that these inequalities hold for all $\theta\in S^{d-1}$. Since ${\hat{\theta}_{\operatorname{relu}}}$ minimizes $\frac{1}{n}\sum_{i=1}^{n}[-\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta]_{+}$, it follows that $\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{\delta}_{{\hat{\theta}_{\operatorname{relu}}}}^{i}\leq 0$. Therefore, together with Lemma 8 to lower-bound $\operatorname{\mathbb{E}}\left[\boldsymbol{\delta}_{\theta}\right]$ for $\theta={\hat{\theta}_{\operatorname{relu}}}$, we have
$\frac{1}{8}\sqrt{\frac{2}{\pi}}\left\lparen 1-\frac{1}{\beta^{2}}\right\rparen\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert^{2}\leq 2C^{\prime}\Delta_{r_{0}}({\hat{\theta}_{\operatorname{relu}}})\sqrt{\frac{\log(1/\delta_{0})}{\beta n}}+2C^{\prime}\Delta_{r_{0}}({\hat{\theta}_{\operatorname{relu}}})^{3/2}\sqrt{\frac{\log(1/\delta_{0})}{n}}\\ +2C^{\prime}\Delta_{r_{0}}({\hat{\theta}_{\operatorname{relu}}})\frac{\log(1/\delta_{0})}{n}+\varepsilon_{0}\Delta_{r_{0}}({\hat{\theta}_{\operatorname{relu}}})^{3/2}\left\lparen\sqrt{\frac{2\ln(1/\delta_{0})}{n}}+2\sqrt{d}\right\rparen.$
Let $r_{0}=\epsilon$, and suppose that $\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert>\epsilon$, in which case we have $\Delta_{r_{0}}({\hat{\theta}_{\operatorname{relu}}})=\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert$. Then, the display above has $\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert$ on both sides of the inequality, and it simplifies to an inequality in $x=\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert^{1/2}$ of the form $x-b\sqrt{x}-c\leq 0$ for some $b,c\geq 0$, which in turn implies $x\leq 1.5(b^{2}+c)$. Therefore
$\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert\leq C^{\prime\prime}\left\lparen\varepsilon_{0}^{2}\left\lparen\frac{\log(1/\delta_{0})}{n}+d\right\rparen+\sqrt{\frac{\log(1/\delta_{0})}{\beta n}}+\frac{\log(1/\delta_{0})}{n}\right\rparen$ )
for some absolute constant $C^{\prime\prime}>0$. (Recall that we have assumed $\beta\geq 1+c$ for some absolute constant $c>0$, and hence $1-1/\beta^{2}\geq c(2+c)/(1+c)^{2}$.) Plug-in $\varepsilon_{0}=\sqrt{\epsilon/(4C^{\prime\prime}d)}$, $\delta_{0}=\delta/(J+\sum_{j=0}^{J}\lvert T_{j}\rvert)$, and
$n\geq\frac{16\log(1/\delta_{0})}{\beta\epsilon^{2}}+\frac{4C^{\prime\prime}\log(1/\delta_{0})}{\epsilon}$
to conclude that $\lVert{\hat{\theta}_{\operatorname{relu}}}-\theta^{\star}\rVert\leq\epsilon$. The lower bound $n$ comes from the assumption in the theorem statement, and the fact that $J=O(\log(1/\epsilon))$ and $\sum_{j=0}^{J}\lvert T_{j}\rvert=O(d/\epsilon)^{O(d)}$. ∎
3.3 Low temperatures
The objective function minimized by ${\hat{\theta}_{\operatorname{relu}}}$ in Equation 2 uses the magnitude of the inner product $\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta$ whenever its sign differs from that of $\mathbf{y}_{i}$. At sufficiently low temperatures ($\beta\gtrsim 1/\epsilon$), this magnitude information can be safely ignored. Specifically, we show that the empirical (zero-one loss) risk minimizer
${\hat{\theta}_{\operatorname{ERM}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})\in\operatorname*{\arg\min}_{\theta\in S^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta\leq 0\}$
achieves near-optimal sample complexity in this regime.
Theorem 4.
Fix any $\epsilon,\delta\in\lparen 0,1\rparen$ ), and assume
$\beta\geq\frac{4\sqrt{2\pi}}{\epsilon}$
and
$n\geq\frac{C(d\log(1/\epsilon)+\log(1/\delta))}{\epsilon}$
where $C>0$ is an absolute constant. Then with probability at least $1-\delta$, we have
$\lVert{\hat{\theta}_{\operatorname{ERM}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert\leq\epsilon.$
Note that the sample size requirement given in Theorem 4 removes a $\log(d)$ factor from that in Theorem 3 in the low temperature regime.
@@ 3.3 Low temperatures
The proof of Theorem 4 is largely based on the following standard performance guarantee for empirical risk minimization, combined with the fact that the VC dimension of the class of homogeneous linear classifiers is $d$ [^2] Note that Lemma 9 is true under every distribution for $(\mathbf{x},\mathbf{y})$; it does not rely on specific properties of our model. In the special case of where $\mathbf{x}$ is the $d$-dimensional standard normal (or any other spherically symmetric distribution) and $\beta=\infty$ (which implies $\operatorname{opt}=0$), the $\log(1/\varepsilon)$ can be removed (Long, 1995). So, in this case, the sample size requirement is $C(d+\log(1/\delta))/\varepsilon$.
Lemma 9 (Vapnik and Chervonenkis, 1971).
@@
$n\geq C\left\lparen\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}\operatorname{opt}+\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right\rparen,$
@@
The following lemma (proved in Appendix C) relates $\lVert{\hat{\theta}_{\operatorname{ERM}}}-\theta^{\star}\rVert$ to $\operatorname{err}_{\theta^{\star}}({\hat{\theta}_{\operatorname{ERM}}})-\operatorname{err}_{\theta^{\star}}(\theta^{\star})$.
@@
For any $\theta\in S^{d-1}$,
$\lVert\theta-\theta^{\star}\rVert\leq\pi\left\lparen\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen+\frac{2\sqrt{2\pi}}{\beta}.$
Proof of Theorem 4.
Note that
Therefore, we can combine Lemma 9 (with $\varepsilon=\epsilon/(2\pi)$) and Lemma 10 with the bound on $\operatorname{opt}=\operatorname{err}_{\theta^{\star}}(\theta^{\star})$ from Lemma 4 to obtain the desired bound on $\lVert{\hat{\theta}_{\operatorname{ERM}}}-\theta^{\star}\rVert$. ∎
3.4 Adaptivity
If the inverse temperature $\beta$ is unknown, then we need to determine which of the aforementioned estimators to use in a data-driven fashion. Notice that the estimators ${\hat{\theta}_{\operatorname{linear}}}$, ${\hat{\theta}_{\operatorname{relu}}}$, and ${\hat{\theta}_{\operatorname{ERM}}}$ may be computed without explicit knowledge of $\beta$. Therefore, it suffices to (coarsely) distinguish between the high ($\beta\lesssim 1$) and moderate-or-low ($\beta\gtrsim 1$) temperature regimes. This can be done using an estimate of $\operatorname{err}_{\theta^{\star}}(\theta^{\star})$ (e.g., training error rate of ${\hat{\theta}_{\operatorname{ERM}}}$) and reasoning about its relationship to $\beta$.